# arXiv:2511.09500v3[stat.ML] 01 Mar 2026

### v3

@@ Abstract
We revisit the problem of denoising from noisy measurements where only the noise level is known, not the noise distribution. In multi-dimensions, independent noise $Z$ corrupts the signal $X$, resulting in the noisy measurement $Y=X+\sigma Z$, where $\sigma\in(0,1)$ is a known noise level. Our goal is to recover the underlying signal distribution $P_{X}$ from denoising $P_{Y}$. We propose and analyze universal denoisers that are agnostic to a wide range of signal and noise distributions. Our distributional denoisers offer order-of-magnitude improvements over the Bayes-optimal denoiser derived from Tweedie’s formula, if the focus is on the entire distribution $P_{X}$ rather than on individual realizations of $X$. Our denoisers shrink $P_{Y}$ toward $P_{X}$ optimally, achieving $O(\sigma^{4})$ and $O(\sigma^{6})$ accuracy in matching generalized moments and density functions. Inspired by optimal transport theory, the proposed denoisers are optimal in approximating the Monge-Ampère equation with higher-order accuracy, and can be implemented efficiently via score matching.
Let $q$ represent the density of $P_{Y}$; for optimal distributional denoising, we recommend replacing the Bayes-optimal denoiser,
@@
with denoisers exhibiting less aggressive distributional shrinkage,
@@
Keywords— Universal denoisers, denoising equations, distributional shrinkage, optimal transport, Monge-Ampère equation, generalized moment matching.
@@ 1 Introduction
$\mathbf{T}^{*}(y)=y+\sigma^{2}\nabla\log q(y)\;,\qquad\text{where $q:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is density function of $P_{Y}$}\;,$
@@
Data-level Shrinkage.
@@
Therefore, at the data level, the shrinkage estimator improves the MSE by only a constant factor, not by its order of magnitude in $\sigma^{2}$. In fact the improvement in MSE is a constant ratio $\tfrac{\sigma^{2}(d-\sigma^{2}\mathop{\mathbb{E}}\|\log q(Y)\|^{2})}{\sigma^{2}d}\in(0,1)$ with small $\sigma$. In this small-noise regime, we shall see shortly that taking a distribution-level viewpoint motivates denoisers that improve denoising performance by an order of magnitude.
@@
Distribution-level Shrinkage.
@@
We use standard notations for gradients and Hessians: for a scalar function $\mathbb{R}^{d}\rightarrow\mathbb{R}, f\nabla f$ is the gradient vector in $\mathbb{R}^{d}$, $\nabla^{2}f$ is the Hessian matrix in $\mathbb{R}^{d\times d}$, and $\Delta f=\mathrm{trace}(\nabla^{2}f)$ is the Laplacian operator. For a vector-valued function $\mathbf{T}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, $\nabla\mathbf{T}$ is the Jacobian matrix in $\mathbb{R}^{d\times d}$, and $\det(\nabla\mathbf{T})$ is its determinant. For two symmetric matrices $A,B\in\mathbb{R}^{d\times d}$, we write $A\preceq B$ if BAB-A is positive semi-definite. $I_{d}\in\mathbb{R}^{d\times d}$ is reserved for the identity matrix. For a vector $x$ and a matrix $A$, we use $\|x\|_{\infty}:=\max_{i}|x_{i}|$ and $\|A\|_{\infty}=\max_{i,j}|A_{ij}|$, and we use $\|x\|$ and $\|A\|$ for its Euclidean norm and Frobenius norm, respectively. $\langle x,y\rangle$ and $\langle A,B\rangle$ denote the Euclidean inner product between two vectors and the Frobenius inner product between two matrices, respectively. We use standard big-O notation: for two functions $f,g:\mathbb{[}0,1]\rightarrow\mathbb{R}$, we write $f(\eta)=O(g(\eta))$ if there exists a constant $C>0$ such that $|f(\eta)/g(\eta)|\leq C$ for all sufficiently small $\eta>0$.
Organizations
@@ 2 Relations to the Literature
Optimal transport theory [Vil08, AGS05] offers a new framework for comparing and transforming probability distributions. Taking an optimal transport perspective also opens new avenues for revisiting the denoising problem. Due to the variational formulation of the Fokker–Planck equation [JKO98], adding small Gaussian noise to a distribution can be interpreted as gradient descent (with a small step size) in the space of probability measures with the entropy functional as the objective and Wasserstein distance as the geometry. Therefore, denoising can be viewed as the reverse process of this gradient descent; see Proposition 1 in [LDK26]. As a result, we can study denoising from the perspective of transport maps. Recent works have explored the use of optimal transport maps for denoising [Kim+23, Che+23, LDK26] and generative modeling [HGL24, LDK26, DL25]. Closest to ours are [GS24, JIS25, Lia25]. [GS24, JIS25] also notice the over-shrinkage effect of the empirical Bayes denoiser. They suggest post-processing empirical Bayes methods to correct for distributional matching, assuming knowledge of $P_{X}$ or its moments. In contrast, our denoisers are universal and do not require any knowledge of $P_{X}$. [Lia25] studies distributional shrinkage under Gaussian noise in a one-dimensional setting but derives a hierarchy of denoisers that leverage higher-order score functions. In contrast, our work studies the problem in a multi-dimensional setting and constructs first- and second-order denoisers that remain universal to a broad range of noise distributions beyond Gaussian.
@@ 3.2 Assumptions on the Signal and Noise
Assumption 1 ($k$-th Order Smoothness).
@@
The second assumption is made on the noise variable $Z\in\mathbb{R}^{d}$. We require certain moment conditions on $Z$. In particular, we do not assume Gaussianity nor independence across coordinates for $Z$. The first-order denoiser defined in (3.1) remains universal and agnostic to the exact noise distribution under Assumption 2(i), whereas the second-order denoiser defined in (3.2) requires both Assumption 2(i) & (ii).
Assumption 2 (Moments of Noise).
@@
As we shall see, for the theoretical result on the first-order denoiser, only Assumption 2(i) and Assumption 1 up to $k=4$, are needed. In particular, we only require the noise variable $Z$ to be uncorrelated, $\mathop{\mathbb{E}}[ZZ^{\top}]=I_{d}$, and to have bounded 4-th moments; no need to assume Gaussianity or independence across coordinates. In this sense, the first-order denoiser is universal to a large class of noise distributions.
For the theoretical result on the second-order denoiser, both Assumption 2(i) & (ii) are needed, and Assumption 1 up to $k=6$. Admittedly, Assumption 2(ii) is more restrictive, as it requires the noise $Z$ to have Gaussian-like $4$-th moments. However, we note that this assumption remains significantly weaker than assuming Gaussianity. In this context, the second-order denoiser is less universal than the first-order denoiser but still agnostic to the specific form of noise distribution. The fourth-moment condition can be relaxed to allow for arbitrary $Z$ with i.i.d. coordinates with a finite fourth-moment, at the cost of a slightly more complex denoising equation, which will be discussed in Section 5.
@@ 4 Matching Generalized Moments
Consider a host of test functions $m:\mathbb{R}^{d}\rightarrow\mathbb{R}\in\mathcal{M}$, we compare how the moments of $\mathbf{T}\sharp P_{Y}$ and $P_{X}$ match:
@@
Assumption 3 (Integrability of Denoisers).
@@
[…]
@@ 4.1 First-Order Denoiser
Definition 1 (2-Smooth Test Functions).
@@
[…]
@@
Theorem 1 (Moment Matching: Optimal First-Order Denoiser).
Under Assumption 3(i), we consider the first-order denoiser $\mathbf{T}_{1}^{f_{*}}(y)$ defined in (3.1),
@@
Let $p=P_{X}$ satisfy Assumption 1 for up to $k=4$, and the noise $Z$ has bounded 4-th moments as in Assumption 2(i). We then have for $\forall m\in\mathcal{M}^{2}(\mathbb{R}^{d})$, the $2$-smooth test function as in Definition 1,
@@
It is possible to state this theorem uniformly over a class of test functions $\mathcal{M}^{2}(\mathbb{R}^{d})$, by imposing uniform bounds on the derivatives of $m\in\mathcal{M}^{2}(\mathbb{R}^{d})$, and uniform integrability conditions. For simplicity, we present the result for each fixed test function $m$. This is because to apply the result to an unbounded test function $m$, one may need to look at a sequence of bounded approximations of $m$, as discussed in the explanations after Definition 1. Such an approximation will require the uniform bounds on the derivatives to grow at a specific rate, which in turn affects the constant $C_{m}$. This is not the focus of this paper, and we leave it to future work.
@@
Corollary 1 (Moment Matching: Bayes-optimal Denoiser).
Consider the same setting and assumptions as in Theorem 1. The Bayes-optimal denoiser
@@
When both the signal and noise are Gaussian distributions, one can take $m=-\log q$, a quadratic function. Symbolically, the leading term on the right-hand side of Corollary 1 is negative, confirming the intuition of over-shrinkage in second moments by the Bayes-optimal denoiser.
@@ 4.2 Second-Order Denoiser
Definition 2 (3-Smooth Test Functions).
@@
[…]
Theorem 2 (Moment Matching: Optimal Second-Order Denoiser).
Under Assumption 3(ii), we consider the second-order denoiser $\mathbf{T}_{2}^{f_{*},g_{*}}$ defined in (3.2),
@@
Let $p=P_{X}$ satisfy Assumption 1 for up to $k=6$, and the noise $Z$ has bounded 6-th moments as in Assumption 2(i) & (ii). We then have for $\forall m\in\mathcal{M}^{3}(\mathbb{R}^{d})$, the $3$-smooth test function as in Definition 2,
@@
This theorem requires slightly stronger conditions on the signal and noise distributions compared to Theorem 1. In particular, the signal distribution $P_{X}$ needs to be smooth up to 6-th order, and the noise distribution $P_{Z}$ needs to have bounded 6-th moments and a Gaussian-like $4$-th moment, namely $\mathop{\mathbb{E}}[Z_{i}^{4}]=3$. Extensions to generic $\mathop{\mathbb{E}}[Z_{i}^{4}]=\kappa\in[1,\infty)$ for noise distribution with i.i.d. coordinates are possible, at the price of a slightly more complex second-order denoising equation. The result holds without requiring Gaussian assumptions, merely the knowledge of the 4-th moments of the noise. Under these conditions, the second-order denoiser $\mathbf{T}_{2}^{f_{*},g_{*}}$ matches the generalized moments (induced by $m$) of the signal distribution $P_{X}$ up to an error of order $O(\sigma^{6})$. This is an order-of-magnitude improvement over the first-order denoiser $\mathbf{T}_{1}^{f_{*}}$ that only achieves $O(\sigma^{4})$ accuracy.
@@ 5 Optimal Denoising and Monge-Ampère Equations
This affinity stems from the study of static Monge-Ampère equations for optimal transport maps. It directly compares the pointwise match between the densities. Suppose $\mathbf{T}$ is the optimal transport map from $P_{Y}$ to $P_{X}$, then the static Monge-Ampère equation [Caf92] states that
@@
In the case when $\mathbf{T}:\Omega\rightarrow\Omega_{X}$ is a diffeomorphism on the support $\Omega\subseteq\mathbb{R}^{d}$ with $\inf_{y\in\Omega}\det(\nabla\mathbf{T}(y))>0$, then $d_{\mathrm{MA}}(\mathbf{T}\sharp P_{Y},P_{X})=0$ if and only if $d_{\mathrm{TV}}(\mathbf{T}\sharp P_{Y},P_{X})=0$. In fact, the MA affinity directly quantifies how the densities match pointwise $|p_{\mathbf{T}(Y)}(x)-p_{X}(x)|$, after accounting for the volume change induced by the transport map $x=\mathbf{T}(y)$.
@@
Assumption 4 (Small Noise Level).
@@
As we evaluate the densities directly in this section, we require the following boundedness assumption on the denoisers induced by $f_{*},g_{*}$ over $\Omega$, the domain of interest. This regularity condition is stronger than Assumption 3 and requires pointwise boundedness. For simplicity, we only concern a compact domain $\Omega\subseteq\mathbb{R}^{d}$, which may be taken as a subset of the support of $P_{Y}$.
Assumption 5 (Boundedness of Denoisers).

### v4

# Distributional Shrinkage I: Universal Denoiser

@@ Abstract
We study the problem of denoising when only the noise level is known, not the noise distribution. Independent noise $Z$ corrupts a signal $X$, yielding the observation $Y=X+\sigma Z$ with known $\sigma\in(0,1)$. We propose universal denoisers, agnostic to both signal and noise distributions, that recover the signal distribution $P_{X}$ from $P_{Y}$. When the focus is on distributional recovery of $P_{X}$ rather than on individual realizations of $X$, our denoisers achieve order-of-magnitude improvements over the Bayes-optimal denoiser derived from Tweedie’s formula, which achieves $O(\sigma^{2})$ accuracy. They shrink $P_{Y}$ toward $P_{X}$ with $O(\sigma^{4})$ and $O(\sigma^{6})$ accuracy in matching generalized moments and densities. Drawing on optimal transport theory, our denoisers approximate the Monge–Ampère equation with higher-order accuracy and can be implemented efficiently via score matching.
Let $q$ denote the density of $P_{Y}$. For distributional denoising, we propose replacing the Bayes-optimal denoiser,
@@
with denoisers exhibiting less-aggressive distributional shrinkage,
@@
Keywords— universal denoisers, generalized moment matching, distributional shrinkage, deconvolution, optimal transport, Monge-Ampère equation, Tweedie’s formula.
@@ 1 Introduction
$\mathbf{T}^{*}(y)=y+\sigma^{2}\nabla\log q(y)\;,\qquad\text{where $q:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the density function of $P_{Y}$}\;,$
@@
Data-level Shrinkage
@@
Therefore, at the data level, the shrinkage estimator improves the MSE by only a constant factor, not by its order of magnitude in $\sigma^{2}$. In fact, the improvement in MSE is a constant ratio $\tfrac{\sigma^{2}(d-\sigma^{2}\mathop{\mathbb{E}}\|\log q(Y)\|^{2})}{\sigma^{2}d}\in(0,1)$ with small $\sigma$. In this small-noise regime, we shall see shortly that taking a distribution-level viewpoint motivates denoisers that improve denoising performance by an order of magnitude.
@@
Distribution-level Shrinkage
@@
We use standard notations for gradients and Hessians: for a scalar function $\mathbb{R}^{d}\rightarrow\mathbb{R}, f\nabla f$ is the gradient vector in $\mathbb{R}^{d}$, $\nabla^{2}f$ is the Hessian matrix in $\mathbb{R}^{d\times d}$, and $\Delta f=\mathrm{trace}(\nabla^{2}f)$ is the Laplacian operator. For a vector-valued function $\mathbf{T}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, $\nabla\mathbf{T}$ is the Jacobian matrix in $\mathbb{R}^{d\times d}$, and $\det(\nabla\mathbf{T})$ is its determinant. For two symmetric matrices $A,B\in\mathbb{R}^{d\times d}$, we write $A\preceq B$ if BAB-A is positive semi-definite. $I_{d}\in\mathbb{R}^{d\times d}$ is reserved for the identity matrix. For a vector $x$ and a matrix $A$, we use $\|x\|_{\infty}:=\max_{i}|x_{i}|$ and $\|A\|_{\infty}=\max_{i,j}|A_{ij}|$, and we use $\|x\|$ and $\|A\|$ for its Euclidean norm and Frobenius norm, respectively. $\langle x,y\rangle$ and $\langle A,B\rangle$ denote the Euclidean inner product between two vectors and the Frobenius inner product between two matrices, respectively. We use standard big-O notation: for two functions $f,g:[0,1]\rightarrow\mathbb{R}$, we write $f(\eta)=O(g(\eta))$ if there exists a constant $C>0$ such that $|f(\eta)/g(\eta)|\leq C$ for all sufficiently small $\eta>0$.
Organization
@@ 2 Relations to the Literature
Optimal transport theory [Vil08, AGS05] offers a new framework for comparing and transforming probability distributions. Taking an optimal transport perspective also opens new avenues for revisiting the denoising problem. Due to the variational formulation of the Fokker–Planck equation [JKO98], adding small Gaussian noise to a distribution can be interpreted as gradient descent (with a small step size) in the space of probability measures with the entropy functional as the objective and Wasserstein distance as the geometry. Therefore, denoising can be viewed as the reverse process of this gradient descent; see Proposition 1 in [LDK26]. As a result, we can study denoising from the perspective of transport maps. Recent works have explored the use of optimal transport maps for denoising [Kim+23, Che+23, LDK26] and generative modeling [HGL24, LDK26, DL25]. Closest to ours are [GS24, JIS25, Lia25]. [GS24, JIS25] also notice the over-shrinkage effect of the empirical Bayes denoiser. They suggest post-processing empirical Bayes methods to correct for distributional matching, assuming knowledge of $P_{X}$ or its moments. In contrast, our denoisers are universal and do not require any knowledge of $P_{X}$. [Lia25] studies distributional shrinkage under Gaussian noise in a one-dimensional setting but derives a hierarchy of denoisers that leverage higher-order score functions. In contrast, our work studies the problem in a multidimensional setting and constructs first- and second-order denoisers that remain universal to a broad range of noise distributions beyond Gaussian.
@@ 3.2 Assumptions on the Signal and Noise
Assumption 3.1 ($k$-th Order Smoothness).
@@
The second assumption is made on the noise variable $Z\in\mathbb{R}^{d}$. We require certain moment conditions on $Z$. In particular, we do not assume Gaussianity nor independence across coordinates for $Z$. The first-order denoiser defined in (3.1) remains universal and agnostic to the exact noise distribution under Assumption 3.2(i), whereas the second-order denoiser defined in (3.2) requires both Assumption 3.2(i) & (ii).
Assumption 3.2 (Moments of Noise).
@@
As we shall see, for the theoretical result on the first-order denoiser, only Assumption 3.2(i) and Assumption 3.1 up to $k=4$, are needed. In particular, we only require the noise variable $Z$ to be uncorrelated, $\mathop{\mathbb{E}}[ZZ^{\top}]=I_{d}$, and to have bounded 4-th moments; no need to assume Gaussianity or independence across coordinates. In this sense, the first-order denoiser is universal to a large class of noise distributions.
For the theoretical result on the second-order denoiser, both Assumption 3.2(i) & (ii) are needed, and Assumption 3.1 up to $k=6$. Admittedly, Assumption 3.2(ii) is more restrictive, as it requires the noise $Z$ to have Gaussian-like $4$-th moments. However, we note that this assumption remains significantly weaker than assuming Gaussianity. In this context, the second-order denoiser is less universal than the first-order denoiser but still agnostic to the specific form of noise distribution. The fourth-moment condition can be relaxed to allow for arbitrary $Z$ with i.i.d. coordinates with a finite fourth-moment, at the cost of a slightly more complex denoising equation, which will be discussed in Section 5.
@@ 4 Matching Generalized Moments
Considering a host of test functions $m:\mathbb{R}^{d}\rightarrow\mathbb{R}\in\mathcal{M}$, we compare how the moments of $\mathbf{T}\sharp P_{Y}$ and $P_{X}$ match:
@@
Assumption 4.1 (Integrability of Denoisers).
@@
(i) $\int_{\mathbb{R}^{d}}\|\nabla f_{*}(y)\|^{2}q(y)\,\mathrm{d}y<\infty$;
(ii) $\int_{\mathbb{R}^{d}}\|\nabla f_{*}(y)\|^{3}+\|\nabla g_{*}(y)\|^{3}q(y)\,\mathrm{d}y<\infty$, and $\sup_{y\in\mathbb{R}^{d}}\|\nabla f_{*}(y)\|^{2}q(y)<\infty$.
@@ 4.1 First-Order Denoiser
Definition 4.2 (2-Smooth Test Functions).
@@
$\mathcal{M}^{2}(\mathbb{R}^{d}):=\left\{m:\mathbb{R}^{d}\rightarrow\mathbb{R}\;\middle|\;\begin{aligned} &\sup_{y\in\mathbb{R}^{d}}\|\nabla^{2}m(y)\|_{\infty}<\infty,\qquad m~\text{vanishes at infinity}\;,\\ &\int_{\mathbb{R}^{d}}|m(y)|\,\mathrm{d}y<\infty,\qquad\int_{\mathbb{R}^{d}}\|\nabla m(y)\|_{\infty}\,\mathrm{d}y<\infty\;.\end{aligned}\right\}\;.$
@@
Theorem 4.3 (Moment Matching: Optimal First-Order Denoiser).
Under Assumption 4.1(i), we consider the first-order denoiser $\mathbf{T}_{1}^{f_{*}}(y)$ defined in (3.1),
@@
Let $p=P_{X}$ satisfy Assumption 3.1 for up to $k=4$, and the noise $Z$ has bounded 4-th moments as in Assumption 3.2(i). We then have for $\forall m\in\mathcal{M}^{2}(\mathbb{R}^{d})$, the $2$-smooth test function as in Definition 4.2,
@@
It is possible to state this theorem uniformly over a class of test functions $\mathcal{M}^{2}(\mathbb{R}^{d})$, by imposing uniform bounds on the derivatives of $m\in\mathcal{M}^{2}(\mathbb{R}^{d})$, and uniform integrability conditions. For simplicity, we present the result for each fixed test function $m$. This is because to apply the result to an unbounded test function $m$, one may need to look at a sequence of bounded approximations of $m$, as discussed in the explanations after Definition 4.2. Such an approximation will require the uniform bounds on the derivatives to grow at a specific rate, which in turn affects the constant $C_{m}$. This is not the focus of this paper, and we leave it to future work.
@@
Corollary 4.4 (Moment Matching: Bayes-optimal Denoiser).
Consider the same setting and assumptions as in Theorem 4.3. The Bayes-optimal denoiser
@@
When both the signal and noise are Gaussian distributions, one can take $m=-\log q$, a quadratic function. Symbolically, the leading term on the right-hand side of Corollary 4.4 is negative, confirming the intuition of over-shrinkage in second moments by the Bayes-optimal denoiser.
@@ 4.2 Second-Order Denoiser
Definition 4.5 (3-Smooth Test Functions).
@@
$\mathcal{M}^{3}(\mathbb{R}^{d}):=\left\{m:\mathbb{R}^{d}\rightarrow\mathbb{R}\;\middle|\;\begin{aligned} &\sup_{y\in\mathbb{R}^{d}}\|\nabla^{3}m(y)\|_{\infty}<\infty,\qquad m,\nabla m~\text{vanish at infinity}\;,\\ &\int_{\mathbb{R}^{d}}|m(y)|\,\mathrm{d}y<\infty,\qquad\int_{\mathbb{R}^{d}}\|\nabla m(y)\|_{\infty}\,\mathrm{d}y<\infty\;.\end{aligned}\right\}\;.$
Theorem 4.6 (Moment Matching: Optimal Second-Order Denoiser).
Under Assumption 4.1(ii), we consider the second-order denoiser $\mathbf{T}_{2}^{f_{*},g_{*}}$ defined in (3.2),
@@
Let $p=P_{X}$ satisfy Assumption 3.1 for up to $k=6$, and the noise $Z$ has bounded 6-th moments as in Assumption 3.2(i) & (ii). We then have for $\forall m\in\mathcal{M}^{3}(\mathbb{R}^{d})$, the $3$-smooth test function as in Definition 4.5,
@@
This theorem requires slightly stronger conditions on the signal and noise distributions compared to Theorem 4.3. In particular, the signal distribution $P_{X}$ needs to be smooth up to 6-th order, and the noise distribution $P_{Z}$ needs to have bounded 6-th moments and a Gaussian-like $4$-th moment, namely $\mathop{\mathbb{E}}[Z_{i}^{4}]=3$. Extensions to generic $\mathop{\mathbb{E}}[Z_{i}^{4}]=\kappa\in[1,\infty)$ for noise distribution with i.i.d. coordinates are possible, at the price of a slightly more complex second-order denoising equation. The result holds without requiring Gaussian assumptions, merely the knowledge of the 4-th moments of the noise. Under these conditions, the second-order denoiser $\mathbf{T}_{2}^{f_{*},g_{*}}$ matches the generalized moments (induced by $m$) of the signal distribution $P_{X}$ up to an error of order $O(\sigma^{6})$. This is an order-of-magnitude improvement over the first-order denoiser $\mathbf{T}_{1}^{f_{*}}$ that only achieves $O(\sigma^{4})$ accuracy.
@@ 5 Optimal Denoising and Monge-Ampère Equations
This affinity stems from the study of static Monge-Ampère equations for optimal transport maps. It directly compares the point-wise match between the densities. Suppose $\mathbf{T}$ is the optimal transport map from $P_{Y}$ to $P_{X}$, then the static Monge-Ampère equation [Caf92] states that
@@
In the case when $\mathbf{T}:\Omega\rightarrow\Omega_{X}$ is a diffeomorphism on the support $\Omega\subseteq\mathbb{R}^{d}$ with $\inf_{y\in\Omega}\det(\nabla\mathbf{T}(y))>0$, then $d_{\mathrm{MA}}(\mathbf{T}\sharp P_{Y},P_{X})=0$ if and only if $d_{\mathrm{TV}}(\mathbf{T}\sharp P_{Y},P_{X})=0$. In fact, the MA affinity directly quantifies how the densities match point-wise $|p_{\mathbf{T}(Y)}(x)-p_{X}(x)|$, after accounting for the volume change induced by the transport map $x=\mathbf{T}(y)$.
@@
Assumption 5.1 (Small Noise Level).
@@
As we evaluate the densities directly in this section, we require the following boundedness assumption on the denoisers induced by $f_{*},g_{*}$ over $\Omega$, the domain of interest. This regularity condition is stronger than Assumption 4.1 and requires point-wise boundedness. For simplicity, we only concern a compact domain $\Omega\subseteq\mathbb{R}^{d}$, which may be taken as a subset of the support of $P_{Y}$.
Assumption 5.2 (Boundedness of Denoisers).